# customstyle linecolor=gray!100, linewidth=3pt, innerleftmargin=3pt, topline=false, rightline=false, bottomline=false, leftline=true, innerrightmargin=3pt, innertopmargin=3pt, innerbottommargin=3pt, backgroundcolor=gray!15

### v1

@@ Abstract.
Atomic commits, each of which addresses a single development concern, are a best practice in software development. However, developers frequently produce tangled commits that mix unrelated changes due to practical constraints or unclear boundaries, negatively impacting code review and maintenance. Although prior commit untangling approaches—rule-based, feature-based, or graph-based—have made progress, they often rely on shallow signals and fail to distinguish between explicit dependencies (e.g., control/data flow) and implicit ones (e.g., semantic or conceptual relationships). In this paper, we propose ColaUntangle, a new collaborative consultation framework for commit untangling that models both explicit and implicit dependencies among code changes. ColaUntangle integrates Large Language Model (LLM)-driven agents in a multi-agent architecture: one agent specializes in explicit dependencies, another in implicit ones, and a reviewer agent synthesizes their perspectives through iterative consultation. To capture explicit and implicit contextual information, we construct multi-version Program Dependency Graphs ($\delta$-PDG), enabling agents to reason over code relationships with both symbolic and semantic depth. We evaluate ColaUntangle on two widely-used datasets (1,612 C# and 14k Java tangled commits). Experimental results show that ColaUntangle outperforms the best-performing baseline, achieving an improvement of 44% on the C# dataset and 100% on the Java dataset. These findings highlight the potential of LLM-based collaborative frameworks for advancing automated commit untangling tasks.
Commit untangling, Atomic commits, Explicit and implicit dependencies, LLMs, Multi-agent collaboration
@@ 1. Introduction
In collaborative software development, a commit is the basic unit of code change, consisting of source file modifications and a message summarizing their purpose (shen2021smartcommit, ). Ideally, each commit should be atomic, addressing a single concern—such as implementing a feature, fixing a bug, or refactoring (li2022utango, ; guo2019decomposing, ). Atomic commits are a widely endorsed best practice for enhancing readability, facilitating reviews, and aiding maintenance by clearly linking changes to their intent (tao2012empiricalwork, ; rigby2012contemporary, ; wang2019cora, ; herbold2022fine, ). As a result, many software communities and companies recommend small, single-purpose commits (fan2024HD-GNN, ; linuxSubmittingPatches, ; microsoftGitGuidance, ).
@@
However, in practice, developers often create tangled commits, which combine code changes related to multiple development concerns within a single commit (fan2024HD-GNN, ; murphy2011we, ). Tangled commits commonly arise due to time constraints or unclear boundaries between concerns during implementation (fan2024HD-GNN, ; parțachi2020flexeme, ; li2022utango, ). For example, a developer may simultaneously refactor code while fixing a bug alongside documentation updates, resulting in a commit that does not adhere to the principle of separation of concerns (herzig2016impact, ; herzig2013impact, ; nguyen2013filtering, ).
@@
Empirical studies (herzig2013impact, ; herzig2016impact, ; tao2015partitioning, ; nguyen2013filtering, ) find that tangled commits are prevalent, indicating that 11% to 40% of commits in software repositories are tangled. Such commits hinder program comprehension, complicate maintenance (tao2012empiricalwork, ; parțachi2020flexeme, ; li2015gated, ). Moreover, tangled commits reduce the accuracy of automated tools such as bug prediction and localization models that rely on mining repository history (herzig2013impact, ; herzig2016impact, ; zhou2019devign, ; kirinuki2014hey, ). These tools often assume that each commit addresses a single concern, so tangled commits introduce noise and degrade model performance.
@@
Researchers have proposed various approaches to untangling commits to split a tangled commit into separate coherent changes. These methods fall into three categories based on their automation level: heuristic rule-based, feature-based, and graph clustering-based. Heuristic methods use manually defined rules—such as line distance, textual similarity, or co-change frequency—to assess change relatedness (barnett2015Heuristic-rule, ; herzig2013impact, ; kirinuki2016splitting, ; wang2019cora, ). Feature-based methods employ handcrafted features (e.g., same method, class, or package) in supervised models to classify change pairs (dias2015EpiceaUntangler, ; yamashita2020changebeadsthreader, ). Graph clustering-based methods build program graphs (e.g., PDGs) to model structural dependencies and use clustering or node embeddings to group related changes (parțachi2020flexeme, ; li2022utango, ; xu2025COMUNT, ; fan2024HD-GNN, ).
@@
Recent advances in large language models (LLMs) have demonstrated strong performance in software engineering tasks that require semantic understanding, reasoning, and explanation (yang2024harnessing, ; fan2023large, ; hou2024large, ). LLMs are particularly well-suited to identify implicit dependencies that go beyond what is detectable via static analysis or statistical patterns. However, a single LLM acting alone may not consistently balance different dependency perspectives or resolve ambiguous cases. Inspired by recent work on multi-agent collaboration (he2025llm, ; wang2025colacare, ; tang2023medagents, ; Kim2024mdagents, ), we propose decomposing the untangling task into subtasks, each handled by specialized agents, and enabling them to collaborate through consultation.
@@
To this end, we introduce ColaUntangle, a collaborative consultation framework for commit untangling that integrates both explicit and implicit dependencies. We construct multi-version Program Dependency Graphs ($\delta$-PDG) to capture structural and contextual information for code changes. Then, we design a multi-agent architecture driven by LLMs: an explicit worker agent focused on explicit dependencies (e.g., control/data flow), an implicit worker agent focused on implicit dependencies (e.g., conceptual relationships), and a reviewer agent that synthesizes and reconciles their results. All agents provide not only untangling decisions but also explanations. Through iterative interaction, these agents simulate human-like consultation and collectively decide on the final untangling outcome.
@@
To evaluate the effectiveness and efficiency of the proposed approach, we conduct experiments on the widely-used C# and Java datasets (parțachi2020flexeme, ; li2022utango, ; fan2024HD-GNN, ) with 1,612 and 14k tangled commits. The evaluation results show that ColaUntangle achieves better untangling results (i.e., 44% and 100% enhancement of effectiveness for C# and Java, compared to the best-performing baseline). Overall, our work makes the following contributions:
@@
  • ColaUntangle: a collaborative consultation model leveraging LLM-driven agents for commit untangling. We propose ColaUntangle, the first LLM-driven collaborative consultation model that untangles commits by considering explicit and implicit dependencies among code changes, simulating human collaborative decision-making to generate results and explanations.
@@
  • Definition of explicit and implicit dependencies for commit untangling. We define a clear and generalizable set of explicit and implicit dependencies between code changes, providing principled guidance that enables both automated and explainable commit untangling.
  • Extensive empirical evaluation. We evaluate ColaUntangle against the state-of-the-art approaches, and the results demonstrate its superior performance. Our tool and data are available at https://anonymous.4open.science/r/ColaUntangle-0109.
2. Motivation and Key Ideas
2.1. Motivating Examples
The following real-world examples serve to motivate ColaUntangle and to introduce its fundamental concepts.
@@ 2.1. Motivating Examples
Example 1. Figure 1A illustrates the changes made in commit 1ad6dbc of the Elasticsearch project. This atom commit updates the index used in the testMultipleBatchesWithLookupJoin function from "idx" to "no_sort_field_idx". The commit includes three separate code changes: lines 3–4 create the new index, lines 6–9 query using the updated index, and lines 11–12 delete the index. These changes do not exhibit any obvious data or control dependencies, nor do they share structural similarities in their syntax. However, they are semantically related, as they collectively form a complete and independent development activity involving the creation, usage, and deletion of an index. This semantic coherence indicates that they belong to a single concern. Existing untangling approaches may fail to handle such commits because they primarily rely on explicit data dependencies or structural similarities between code changes, overlooking the semantic relationships that often define a developer’s intent.
Observation 1 [Semantic]. The semantics of code changes are crucial for determining their underlying concern. Changes that belong to the same concern often share semantic similarity.
@@
Example 2. Figure 1B shows another atom commit. Lines 3–6 define an enum class representing ordering options, and lines 8–16 implement a sorting method that utilizes this enum (only the if block is shown in the figure, as the else block has a similar structure). The code changes in lines 3–6 and lines 8–16 exhibit logical dependencies respectively, as they belong to the same class or method implementation. Furthermore, there is a clear data dependency: line 11 directly uses the value defined in the enum class introduced in lines 3–6. Due to these logical and data dependencies, these changes are categorized as belonging to a single concern.
Observation 2 [Data and control dependency]. Code changes within the same concern often exhibit data or control dependencies.
@@
Example 3. Figure 1C shows a tangled commit containing two distinct concerns: ”remove LicensingCharset” and ”reformat code”. The first concern (at lines 3 and 6) involves changes to the charset used in the TrialLicenseUtils class (only a subset of these changes is shown, as other charset modifications exhibit similar structure). Each of these changes is identical, replacing LicensesCharset with StandardCharsets. The affected statements are cloned code segments used to support different variables (issueTo and feature), resulting in high similarity in both textual content and code structure.
Observation 3 [Code similarity]. Code changes within the same concern often exhibit high similarity in text and structure.
@@
The second concern focuses on reformatting code at lines 9, and 13, spanning two files. These changes represent typical cosmetic edits, which include syntactic formatting, refactoring, or non-functional textual modifications. Such edits are common in daily software development and usually belong to a single concern. However, existing approaches fail to detect and classify these changes correctly, as their rules, features, or graph representations do not model cosmetic edits.
Observation 4 [Cosmetic edit]. Code changes caused by cosmetic edits, such as syntactic formatting, refactoring, or non-functional textual modifications, often belong to the same concern.
2.2. Key Ideas
Inspired by the above observations, we propose ColaUntangle with the following key ideas.
2.2.1. Key Idea 1 [Explicit and Implicit Dependency]
@@
Based on Observations 1–4, we find that code changes can be grouped into different concerns due to various factors. To provide comprehensive guidance for our method, we categorize the characteristics of code changes belonging to the same concern into two types: explicit dependencies and implicit dependencies.
@@
  • Code changes with data dependencies often belong to the same concern. [Observation 2]
  • Code changes with control dependencies often belong to the same concern. [Observation 2]
@@
  • Code changes with semantic similarity may belong to the same concern. [Observation 1]
  • Code changes with high textual or structure similarity may belong to the same concern. [Observation 3]
  • Code changes introduced for cosmetic edits, such as syntactic formatting, refactoring, or non-functional textual modifications, often belong to the same concern. [Observation 4]
2.2.2. Key Idea 2 [LLM-driven Foundation Models]
Building on Observation 1, 3 and 4, we decide to leverage the powerful capabilities of LLMs rather than traditional machine learning or neural network models to untangle commits. LLMs exhibit superior performance in comprehending code semantics, performing reasoning, and adapting to diverse decision-making tasks (liu2024large, ; jin2024llms, ). By employing LLM-driven agents, ColaUntangle can effectively understand the intent and implicit dependencies underlying code changes beyond purely structural or syntactic similarities. This enables our approach to untangle commits with a deeper semantic understanding which is difficult to achieve based on heuristic rules, feature representations and graph models.
2.2.3. Key Idea 3 [Collaborative Consultation]
As seen in Observations 1–4, code changes may belong to different concerns due to diverse and sometimes conflicting factors, making it difficult to define a single generalized untangling criterion. Therefore, we design a multi-agent collaborative consultation process comprising distinct roles and iterative negotiation. Specifically, we define two types of worker agents: one specializing in analyzing explicit dependencies between code changes and the other focusing on implicit dependencies. Additionally, we introduce a reviewer agent that synthesizes and integrates the outputs from both worker agents. These three agents engage in multiple rounds of collaborative consultation, producing the final untangling result upon reaching consensus.

### v2

# LLM-Driven Collaborative Model for Untangling Commits via Explicit and Implicit Dependency Reasoning

@@ Abstract.
Atomic commits, which address a single development concern, are a best practice in software development. In practice, however, developers often produce tangled commits that mix unrelated changes, complicating code review and maintenance. Prior untangling approaches—rule-based, feature-based, or graph-based—have made progress but typically rely on shallow signals and struggle to distinguish explicit dependencies (e.g., control/data flow) from implicit ones (e.g., semantic or conceptual relationships). In this paper, we propose ColaUntangle, a new collaborative consultation framework for commit untangling that models both explicit and implicit dependencies among code changes. ColaUntangle integrates Large Language Model (LLM)-driven agents in a multi-agent architecture: one agent specializes in explicit dependencies, another in implicit ones, and a reviewer agent synthesizes their perspectives through iterative consultation. To capture structural and contextual information, we construct Explicit and Implicit Contexts, enabling agents to reason over code relationships with both symbolic and semantic depth. We evaluate ColaUntangle on two widely-used datasets (1,612 C# and 14k Java tangled commits). Experimental results show that ColaUntangle outperforms the best-performing baseline, achieving an improvement of 44% on the C# dataset and 82% on the Java dataset. These findings highlight the potential of LLM-based collaborative frameworks for advancing automated commit untangling tasks.
Commit untangling, Explicit and implicit dependencies, LLMs, Multi-agent collaboration
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Automatic programmingccs: Software and its engineering Software configuration management and version control systems
@@ 1. Introduction
In collaborative software development, a commit is the basic unit of code change, consisting of source file modifications and a message summarizing their purpose (shen2021smartcommit; dong2022fira; gomez2015visually). Ideally, each commit should be atomic, addressing a single concern—such as implementing a feature, fixing a bug, or refactoring (li2022utango; guo2019decomposing; ram2018makes; robillard2002concern). Atomic commits are a widely endorsed best practice for improving readability, facilitating reviews, and helping maintenance by clearly linking changes to their intention (tao2012empiricalwork; rigby2012contemporary; wang2019cora; herbold2022fine; baysal2016investigating). As a result, many software communities and companies recommend small, single-purpose commits (fan2024HD-GNN; linuxSubmittingPatches; microsoftGitGuidance).
@@
However, in practice, developers often create tangled commits, which combine code changes related to multiple development concerns within a single commit (fan2024HD-GNN; murphy2011we). Tangled commits commonly arise due to time constraints or unclear boundaries between concerns during implementation (fan2024HD-GNN; parțachi2020flexeme; li2022utango). For example, a developer may simultaneously refactor code while fixing a bug alongside documentation updates, resulting in a commit that does not adhere to the principle of separation of concerns (herzig2016impact; herzig2013impact; nguyen2013filtering).
@@
Empirical studies (herzig2013impact; herzig2016impact; tao2015partitioning; nguyen2013filtering) find that tangled commits are prevalent, indicating that 11% to 40% of commits in software repositories are tangled. Such commits hinder program comprehension, complicate maintenance (tao2012empiricalwork; parțachi2020flexeme; li2015gated). Moreover, tangled commits reduce the accuracy of automated tools such as bug prediction and localization models that rely on mining repository history (herzig2013impact; herzig2016impact; zhou2019devign; kirinuki2014hey; zimmermann2005mining; guo2017interactively). These tools often assume that each commit addresses a single concern, so tangled commits introduce noise and degrade model performance.
@@
Researchers have proposed various approaches to untangling commits to split a tangled commit into separate coherent changes. These methods fall into three categories based on their automation level: heuristic rule-based, feature-based, and graph clustering-based. Heuristic methods use manually defined rules—such as line distance, textual similarity, or co-change frequency—to assess change relatedness (barnett2015Heuristic-rule; herzig2013impact; kirinuki2016splitting; wang2019cora). Feature-based methods employ handcrafted features (e.g., same method, class, or package) in supervised models to classify change pairs (dias2015EpiceaUntangler; yamashita2020changebeadsthreader). Graph clustering-based methods build program graphs, e.g., program dependency graphs (PDGs), to model structural dependencies and use clustering or node embeddings to group related changes (parțachi2020flexeme; li2022utango; xu2025COMUNT; fan2024HD-GNN).
@@
Recent advances in large language models (LLMs) have demonstrated strong performance in software engineering tasks that require semantic understanding, reasoning, and explanation (yang2024harnessing; fan2023large; hou2024large; fan2025exploring; li2024understanding). LLMs are particularly well-suited to identify implicit dependencies that go beyond what is detectable via static analysis or statistical patterns. However, a single LLM acting alone may not consistently balance different dependency perspectives or resolve ambiguous cases. Inspired by recent work on multi-agent collaboration (he2025llm; wang2025colacare; tang2023medagents; Kim2024mdagents), we propose decomposing the untangling task into subtasks, each handled by specialized agents, and enabling them to collaborate through consultation.
@@
To this end, we introduce ColaUntangle, a collaborative consultation framework for commit untangling that integrates both explicit and implicit dependencies. We construct Explicit and Implicit Contexts to capture explicit and implicit dependencies between code changes. Then, we design a multi-agent architecture driven by LLMs: an explicit worker agent focused on explicit dependencies (e.g., control/data flow), an implicit worker agent focused on implicit dependencies (e.g., conceptual relationships), and a reviewer agent that synthesizes and reconciles their results. All agents provide not only untangling decisions but also explanations. Through iterative interaction, these agents simulate human-like consultation and collectively decide on the final untangling outcome.
@@
To evaluate the effectiveness and efficiency of the proposed approach, we conduct experiments on the widely-used C# and Java datasets (parțachi2020flexeme; li2022utango; fan2024HD-GNN) with 1,612 and 14k tangled commits. The evaluation results show that ColaUntangle achieves better untangling results (i.e., 44% and 82% enhancement of effectiveness for C# and Java, compared to the best-performing baseline). Overall, our work makes the following contributions:
@@
  • ColaUntangle: the first collaborative consultation model leveraging LLM-driven agents for commit untangling. We propose ColaUntangle, the first LLM-driven collaborative consultation model that untangles commits by considering explicit and implicit dependencies among code changes.
@@
  • Formal and interpretable evidence framework. We define a clear and generalizable set of Explicit and Implicit Dependencies between code changes. This principled evidence framework provides direct guidance for LLM reasoning, which is foundational for explainable untangling.
  • A unified LLM framework for explainable, human-aligned untangling. We introduce a unified framework leveraging LLM-driven reasoning to infer semantics, robustly balance explicit and implicit dependencies, identify cosmetic edits and generate contextual explanations. Unlike prior black-box untangling models, this collaborative and human-centric architecture achieves explainable and interpretable untangling results.
  • Extensive empirical evaluation. We evaluate ColaUntangle against the state-of-the-art approaches, and the results demonstrate its superior performance. Beyond quantitative results, we conduct a detailed examination of both correctly untangled and misclassified cases, revealing how explainable reasoning contributes to accurate decisions and exposing the boundaries of current LLM-based untangling.
2. Motivation
High-quality software development demands atomic commits, where each commit addresses a single, isolated concern. However, in practice, developers frequently introduce tangled commits containing changes related to multiple independent activities. While prior work has attempted to untangle these commits, they often fall short due to an inability to fully capture the diverse and often subtle relationships between code changes. To define the scope of the problem and motivate the design of ColaUntangle, we analyze real-world commit examples, which crystallize four critical challenges faced by existing untangling methods.
2.1. Motivating Examples and Core Challenges
The examples in Figure 1, demonstrate the inadequacy of existing studies and highlight the necessity for an advanced, semantically-aware approach.
@@ 2.1. Motivating Examples and Core Challenges
Figure 1A (Example 1) illustrates the changes made in commit 1ad6dbc of the Elasticsearch project. This atom commit updates the index used in the testMultipleBatchesWithLookupJoin function from "idx" to "no_sort_field_idx". The commit includes three separate code changes: lines 3–4 create the new index, lines 6–9 query using the updated index, and lines 11–12 delete the index. These changes do not exhibit any obvious data or control dependencies, nor do they share structural similarities in their syntax. However, they are semantically related, as they collectively form a complete and independent development activity involving the creation, usage, and deletion of an index. This semantic coherence indicates that they belong to a single concern. Existing untangling approaches may fail to handle such commits because they primarily rely on explicit data dependencies or structural similarities between code changes (parțachi2020flexeme; li2022utango; fan2024HD-GNN), overlooking the semantic relationships that often define a developer’s intent.
Challenge 1: Semantic Disconnect. Precisely untangling commits requires inferring the underlying semantic purpose and grouping changes that lack explicit structural links but form a cohesive, complete development activity.
@@
Figure 1B (Example 2) shows another atom commit. Lines 3–6 define an enum class representing ordering options, and lines 8–16 implement a sorting method that utilizes this enum (only the if block is shown in the figure, as the else block has a similar structure). The code changes in lines 3–6 and lines 8–16 exhibit logical dependencies respectively, as they belong to the same class or method implementation. Furthermore, there is a clear data dependency: line 11 directly uses the value defined in the enum class introduced in lines 3–6. These changes are categorized as belonging to a single concern based on their logical and data dependencies, which are strong structural indicators for grouping.
@@
However, many concerns are defined by implicit signals, where such strong explicit links are absent. Figure 1C (Example 3.) shows a tangled commit containing two distinct concerns: ”remove LicensingCharset” and ”reformat code”. The first concern (at lines 3 and 6) involves changes to the charset used in the TrialLicenseUtils class (only a subset of these changes is shown, as other charset modifications exhibit similar structure). Each of these changes is identical, replacing LicensesCharset with StandardCharsets. The affected statements are cloned code segments used to support different variables (issueTo and feature), resulting in high textual and structural similarity which is a crucial implicit signal for grouping. The variety of relationships between code changes—from the strong, explicit data flow in Example 2 to the subtle, implicit similarity here—presents an integration dilemma. Existing approaches struggle to meaningfully balance these two dependency types. They rely primarily on explicit dependencies, treating implicit signals (like similarity) as secondary factors or only using them for post-hoc adjustments (e.g., UTango (li2022utango)). This inability to fluidly and contextually integrate the entire spectrum of evidence (explicit and implicit dependency) leads to inaccurate or fragile untangling results.
Challenge 2: Contextual Integration. A unified approach must effectively integrate and weigh diverse forms of evidence—ranging from strong explicit dependencies (data/control flow) to subtle implicit relationships (semantics/similarity)—to accurately determine concern boundaries.
@@
The second concern in Figure 1C (Example 3) focuses on reformatting code at lines 9, and 13, spanning two files. These changes represent typical cosmetic edits, which include syntactic formatting, refactoring, or non-functional textual modifications. Such edits are common in daily software development and usually belong to a single concern. However, existing approaches fail to detect and classify these changes correctly, as their rules (barnett2015Heuristic-rule; herzig2016impact), features (dias2015EpiceaUntangler), or graph representations (parțachi2020flexeme; li2022utango; fan2024HD-GNN) do not model cosmetic edits. Due to this oversight, they often fail to correctly isolate cosmetic edits, which can scatter across files and hide within larger functional concerns.
Challenge 3: Specialized Classification. Existing methods lack the capability to accurately identify and isolate non-functional cosmetic changes, leading to incomplete untangling.
Except for this, existing methods act as a black box, failing to provide the contextual reasoning behind their grouping decisions (parțachi2020flexeme; li2022utango; fan2024HD-GNN). This lack of a clear, contextual rationale is a significant limitation in a task requiring complex contextual judgment. A robust solution should offer interpretability, explaining its reliance on specific evidence (e.g., ”grouped due to semantic completion” or ”separated due to clear data independence”).
Challenge 4: Rationale and Explainability. Existing untangling methods lack the capacity to provide a clear, contextual rationale for their grouping decisions.
2.2. Key Principles
Based on the core challenges identified in Section 2.1, we propose ColaUntangle with an architecture founded on three key principles. These principles are designed to overcome the limitations by integrating deep semantic reasoning with a robust, multi-perspective decision-making process.
2.2.1. Principle 1: LLM-driven Foundation Models
To address Challenge 1 (Semantic Disconnect) and Challenge 3 (Specialized Classification), we decide to leverage the powerful capabilities of LLMs rather than traditional machine learning or neural network models to untangle commits. LLMs exhibit superior performance in comprehending code semantics, performing reasoning, and adapting to diverse decision-making tasks (liu2024large; jin2024llms). By employing LLM-driven agents, ColaUntangle can effectively understand the intent and implicit dependencies underlying code changes beyond purely structural or syntactic similarities. This enables our approach to untangle commits with a deeper semantic understanding and accurately classify non-functional edits which is difficult to achieve based on heuristic rules, feature representations and graph models.
2.2.2. Principle 2: Explicit and Implicit Dependency
@@
Based on the motivating examples, we find that code changes can be grouped into different concerns due to various factors. To effectively guide the LLM’s reasoning, we formalize the diverse observations from our motivating examples into a structured framework of evidence. We categorize the characteristics linking code changes into two distinct, orthogonal types: Explicit Dependencies and Implicit Dependencies.
@@
  • Code changes with data dependencies often belong to the same concern.
  • Code changes with control dependencies often belong to the same concern.
@@
  • Code changes with semantic similarity may belong to the same concern.
  • Code changes with high textual or structure similarity may belong to the same concern.
  • Code changes introduced for cosmetic edits, such as syntactic formatting, refactoring, or non-functional textual modifications, often belong to the same concern.
Unlike prior studies that offered vague or indirect definitions of explicit and implicit dependencies, this formal set of rules provides a direct, comprehensive, and interpretable basis for untangling, enabling the LLM agents to reason directly over these untangling rules. The LLM agents can directly reason over these explicit principles. This design is fundamental to achieving both robust evidence integration (Challenge 2) and generating the clear, contextual rationale required for explainability (Challenge 4).
2.2.3. Principle 3: Collaborative Consultation
Recognizing that untangling requires resolving conflicts between the diverse evidence defined in Principle 2, and to address Challenge 2 (Contextual Integration) and Challenge 4 (Rationale and Explainability), we implement a multi-agent collaborative consultation process comprising distinct roles and iterative negotiation. Specifically, we define two types of worker agents: one specializing in analyzing explicit dependencies between code changes and the other focusing on implicit dependencies. Additionally, we introduce a reviewer agent that synthesizes and integrates the outputs from both worker agents. These three agents engage in multiple rounds of collaborative consultation, producing the final untangling result upon reaching consensus.